# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v4

@@ 5 Conclusion
The fundamental objective of the Universal PS task is to reconstruct illumination-invariant surface normals from a series of multi-light input images. Consequently, the decoder must effectively aggregate contextual information provided by the encoder to derive a lighting-agnostic representation of these surface normals. Existing Universal PS methods predominantly utilize encoder-decoder architectures, wherein encoders typically possess a significantly larger parameter count than their decoder counterparts. Therefore, assuming a relatively consistent decoder design, empowering a more capable encoder to effectively disentangle lighting from normal information and refine these into robust normal-specific features is expected to yield superior normal reconstruction.
Building upon approaches such as UniPS Ikehata (2022) and SDM UniPS Ikehata (2023) that utilize comparable decoder designs, a primary driver for our LINO UniPS is the introduction of an improved encoder. The objective of this encoder is to yield more consistent features, thereby achieving superior results. The direct correlation between greater feature consistency and superior normal reconstruction is not an oversimplification; the decoder’s architecture and capabilities also represent a non-negligible factor in overall performance. Nonetheless, the decoders utilized in UniPS, SDM UniPS and our LINO UniPS are architecturally similar, all adhering to the pixel-sampling paradigm. Consequently, when analyzing the relationship between feature consistency (e.g., as measured by CSIM/SSIM) and normal reconstruction performance in Fig. 1, we group these three methods together to facilitate a more direct comparison of the impact of their respective encoder-derived features.
Fig. A7 presents Principal Component Analysis (PCA) visualizations of features extracted by the encoders of various methods, alongside their corresponding feature similarity metrics (CSIM/SSIM). In the following discussion, we primarily focus our analysis on our Ours w/mlp variant and Uni MS-PS.
@@
Furthermore, We need to figure out that Uni MS-PS exhibits certain practical limitations. (a) Its multi-scale nature leads to considerable inference latency, particularly when processing multiple high-resolution input images (e.g., handling 16 images at 4K resolution can extend to about an hour). In contrast, our LINO UniPS method typically completes inference within tens of seconds for similar inputs (see Tab. 6). (b) While Uni MS-PS can reconstruct detailed surface normals, its reliance on potentially lower-resolution training datasets and its patch-based inference mechanism can lead to a loss of global contextual information, sometimes resulting in reconstructions that are locally detailed but globally inconsistent or erroneous (see Fig. 1).
@@
Despite the commendable performance demonstrated by LINO UniPS, certain limitations nonetheless remain, offering avenues for future research.
Firstly, our incorporation of global attention in the encoder, while aimed at enhancing intra-image feature communication for more effective decoupling of lighting from normal features and successfully improving disentanglement, also introduces computational burden. A key direction for future work is therefore to explore more computationally efficient mechanisms that can achieve comparable decoupling efficacy at a reduced operational cost.
Secondly, a limitation arises in the estimation of surface normals for near-planar objects that exhibit intricate concave and convex surface details, where the performance of LINO UniPS can be suboptimal. We attribute this primarily to a fundamental challenge inherent in the Universal Photometric Stereo (PS) paradigm: the absence of explicit light source parameters makes it difficult for the network to unambiguously determine the precise illumination direction (e.g., to distinguish between light originating from above versus below the surface). Consequently, this ambiguity can frequently, and in some instances, even lead to the estimated surface normals being inverted (see Fig. A8). In future investigations, we plan to explore whether more sophisticated lighting alignment strategies could effectively mitigate this specific issue.
@@
[…]
A1.5.2 Real Data

### v5

@@ Acknowledgements.
Universal PS aims to solve an inverse problem: recovering the intrinsic surface normal field $N$ from a set of observations $\{I_{f}\}_{f=1}^{F}$ formed under varying, unknown illumination conditions $\{\mathbf{L}_{f}\}_{f=1}^{F}$. The imaging process can be formally described as $I_{f}=\mathcal{F}(N,\rho,\mathbf{L}_{f})$, where $\mathcal{F}$ represents the rendering equation and $\rho$ the surface reflectance.
An ideal encoder $E$ seeks to extract a unified feature representation $F_{enc}=E(\{I_{f}\}_{f=1}^{F})$ that is strictly illumination-invariant, effectively marginalizing out the extrinsic variable $\{\mathbf{L}_{f}\}_{f=1}^{F}$. In this physical context, feature similarity (CSIM; SSIM) serves as a direct quantitative metric of this invariance; a low CSIM/SSIM score implies that the feature $F_{enc}$ retains significant dependency on $\{\mathbf{L}_{f}\}_{f=1}^{F}$, indicating a failure to decouple extrinsic illumination from intrinsic geometry.
The decoder $D$ is tasked with learning the mapping $\hat{N}=D(\{\mathbf{Z}_{f}\}_{f=1}^{F})$. When features are not decoupled, the decoder confronts a highly ill-posed problem: it receives highly variable inputs $F_{enc}$ for the exact same physical geometry $N$, which inevitably introduces ambiguity and variance into the estimator, manifesting as higher reconstruction error (MAE).
Therefore, a primary driver for our LINO UniPS is the introduction of an improved encoder $E$. By explicitly employing Light Register Tokens supervised by Light Alignment to physically isolate variant illumination components $\{\mathbf{L}_{f}\}_{f=1}^{F}$, our encoder ensures that the feature representation passed to the decoder remains pure and consistent. This effectively mitigates the ill-posed nature of the decoding task, naturally yielding higher accuracy.
While the decoder’s capabilities represent a non-negligible factor in overall performance, the decoders utilized in UniPS, SDM UniPS, and our LINO UniPS are architecturally similar, all adhering to the pixel-sampling paradigm. Consequently, the direct correlation between greater feature consistency and superior normal reconstruction is not an oversimplification; thus, when analyzing this relationship in Fig. 1, we group these three methods together to facilitate a direct comparison of the impact of their respective encoder-derived features.
Fig. A8 presents Principal Component Analysis (PCA) visualizations of features extracted by the encoders of various methods, alongside their corresponding feature similarity metrics (CSIM/SSIM). In the following discussion, we focus our analysis on our Ours w/mlp variant and Uni MS-PS.
@@
Furthermore, we need to figure out that Uni MS-PS exhibits certain practical limitations. (a) Its multi-scale nature leads to considerable inference latency, particularly when processing multiple high-resolution input images (e.g., handling 16 images at 4K resolution can extend to about an hour). In contrast, our LINO UniPS method typically completes inference within tens of seconds for similar inputs (see Tab. 6). (b) While Uni MS-PS can reconstruct detailed surface normals, its reliance on potentially lower-resolution training datasets and its patch-based inference mechanism can lead to a loss of global contextual information, sometimes resulting in reconstructions that are locally detailed but globally inconsistent or erroneous (see Fig. 1 and Fig. 7) .
@@
Despite the commendable performance demonstrated by LINO UniPS, certain limitations remain, offering avenues for future research.
Firstly, the incorporation of global attention within our encoder, while designed to enhance inter-image feature interaction for more effective illumination-normal decoupling and successfully improving disentanglement, inevitably introduces additional computational burden. Consequently, a key direction for future work is to explore more computationally efficient mechanisms that can achieve comparable decoupling efficacy at a reduced operational cost.
Secondly, despite its strong generalization capabilities, it is important to acknowledge that the Universal PS paradigm exhibits certain inherent drawbacks compared to traditional paradigms.
For example, traditional Calibrated PS methods benefit from explicit, known light source parameters. In contrast, Universal PS lacks any explicit light source input (Although we employ Light Register Tokens to mitigate this, our network operates without explicit light source input during inference. However, Calibrated PS methods do utilize). Consequently, when handling near-planar objects, the Universal PS method struggles to unambiguously distinguish whether light originates from "above" or "below", resulting in the reconstructed surface normals being inverted (see Fig. A9), whereas Calibrated PS handles this relatively well.
Furthermore, the Universal PS method is data-driven, inherently relying on massive amounts of data. Conversely, traditional Calibrated and Uncalibrated PS approaches do not rely as heavily on large-scale training datasets.
@@
A1.5.2 Real Data & Synthetic Benchmark
@@
Figure A14: Qualitative results of the effect of each Light Register Token type. The full three-token model (right) achieves higher visual fidelity than any single-token variant, highlighting the need for all three.
A1.6 Additional Ablations
A1.6.1 Effect of Each Light Register Token Type
Table A2: Quantitative comparison of single-token (Direction, Env, Point) performance against the full three-token combination, highlighting the Point token’s contribution and the critical synergy of all three.
[…]
Direction 0.84 0.81 5.32
Env 0.83 0.82 5.30
Point 0.86 0.84 4.98
Direction & Env & Point 0.88 0.86 4.51
To analyze the independent contributions of each proposed Light Register Token, we designed a fine-grained ablation study. In this experiment, we use the final model presented in the last row of Tab. 2 as our baseline. The variable is the specific Light Register Token used, along with their corresponding light alignment losses. We evaluated the following four setups: (a) Using Direction Tokens, (b) Using Env Tokens, (c) Using Point Tokens, (d) Using Point & Direction & Env Tokens (final model). The total number of register tokens was kept at three for all setups (consistent with the final model). All other experimental settings were identical.
Quantitative results are shown in Tab. A2, and qualitative results are in Fig. A14. We observe that: among all single-token configurations, using only the Point Tokens yields the most significant performance improvement, while the effects of the Direction and Env tokens are similar. This is because the Point Tokens represent high-frequency illumination information (such as specular highlights) , which provides the most critical and informative cues for resolving surface normals Ikehata (2023; 2022). Conversely, the Direction and Env tokens correspond to low-frequency light sources, providing relatively fewer geometric cues.
Furthermore, we observe that a significant performance gap still exists between any single-token variant (configurations (a), (b), (c)) and the full model using all three tokens (configuration (d)). This strongly demonstrates that our design is not redundant: real-world illumination is a complex mixture, and the synergistic effect of all three tokens enables the model to handle more complex lighting conditions, thus most thoroughly decoupling the complex illumination information from the normal features.
A1.6.2 Why Specialized Tokens Instead of Generic Tokens
Table A3: Quantitative comparison of Specialized vs. Generic Light Register Tokens. Our specialized token design outperforms the generic token method.
[…]
Generic 0.84 0.84 5.07
Specialized 0.88 0.86 4.51
We introduce three specialized Light Register Tokens to correspond to three distinct illumination types (Point, Direction, and Env). A natural question arises: why not use generic tokens instead of this specialized design? To answer this, we conduct an ablation study. For this experiment, we use our final model (last row of Tab. 2) as the baseline. All other experimental settings remain identical. The only change we make is to replace our three specialized Light Register Tokens with three generic tokens. These generic tokens are simultaneously supervised by all three light alignment losses ($\mathcal{L}_{\text{env}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{direction}}$)
Quantitative results are shown in Tab. A3, and qualitative results are in Fig. A15. We find that the performance of the generic token approach is lower than our specialized token approach. We analyze that this is because the generic tokens are forced to entangle physically disparate lighting cues (e.g., high-frequency and low-frequency light), which consequently degrades the accuracy of the subsequent normal prediction.
Figure A15: Qualitative comparison of Specialized vs. Generic Light Register Tokens. Our specialized token approach (right) exhibits superior visual fidelity.
Figure A16: Qualitative comparison of Dual-branch model against its single-branch variants, confirming the complementary effect.
A1.6.3 Why Using Dual-branch Architecture
This section analyzes the advantages of our Dual-branch architecture (parallel wavelet and naive downsample branches). Quantitative results are in Tab. A4, and qualitative results are in Fig. A16. For this experiment, we use our final model (last row of Tab. 2) as the baseline. All other settings are identical; we only modify the encoder’s feature extraction architecture to create three variants: (1) Downsample-only (2) Wavelet-only (3) Dual-branch (final model).
From both quantitative and qualitative results, we find that Wavelet-only outperforms Downsample-only. For normal reconstruction, naive downsample irreversibly discards critical high-frequency geometric information. In contrast, the wavelet transform, by explicitly preserving these high-frequency components , captures finer surface details and yields performance gain.
Table A4: Quantitative comparison of Dual-branch model against its single-branch variants, demonstrating the quantitative advantage of the Dual-branch design.
[…]
Downsample-only 0.80 0.79 5.98
Wavelet-only 0.85 0.84 4.97
Dual-branch 0.88 0.86 4.51
Furthermore, Dual-branch model outperforms all single-branch variants. Our analysis is as follows: The two branches achieve functional specialization. The wavelet branch focuses on explicitly preserving high-frequency, fine-grained local geometric details. The naive downsample branch, while lossy in detail, provides a smoother, more anti-aliased low-frequency representation, offering robust global image-domain semantic information Glasner et al. (2009); Zhang (2019). The advantages of both branches are complementary Guo et al. (2017), and their fusion allows the model to achieve superior accuracy; therefore, despite the low-frequency component already included in the wavelet transform, we additionally retain the parallel naive downsample branch.